OpenStudy (anonymous):

(x + 8)(x – 6) if you multiply this,would it be x^2-48?

OpenStudy (hoblos):

(a+b)(c+d) = ac+ad+bc+bd (x + 8)(x – 6) = x^2 -6x + 8x - 48 = x^2 +2x -48

OpenStudy (anonymous):

oh oka...thanks

OpenStudy (hoblos):

my pleasure XD